A particle of mass m has speed v = c/x, where c is a constant and x is the displacement. At t=0, x=

x Find the force

a) F(x)

b) F(t)



1
Expert's answer
2021-11-16T13:41:03-0500

"F=ma=m\dot v=-\frac{mc}{x^2},"

a)

"F=-\frac{mc}{t^2},"

b)

"F=-\frac{mc}{x^2}."